A certain molecule has the fundamental and the first overtone centered at 2143.26cm–¹ and 4260.04cm–¹. Determine the vibrational frequency and first anharmonicity constant


1
Expert's answer
2021-02-16T07:46:45-0500

The energy levels for the Morse potential are: 

G(v) = (v + ½)ωe - (v + ½)2ωexe(in cm-1

The fundamental corresponds to the transition between v = 0 and v = 1. This occurs at: 

G(1) - G(0) = [(1 + ½)ωe - (1 + ½)2ωexe] – [(0 + ½)ωe - (0 + ½)2ωexe

 = [(3/2)ωe– (9/4)ωexe] – [(1/2)ωe -(1/4)ωeexe] = ωe – 2ωexe

The first overtone corresponds to the transition between v = 0 and v = 2. This occurs at: 

G(2) - G(0) = [(2 + ½)ωe - (2 + ½)2ωexe] – [(0 + ½)ωe - (0 + ½)2ωexe

 = [(5/2)ωe – (25/4)ωexe] – [(1/2)ωe – (1/4)ωexe] = 2ωe – 6ωexe

As these occur at 3034 cm-1 and 5941 cm-1 respectively; 

ωe – 2ωexe = 2143.26 cm-1 (1) 

e – 6ωexe = 4260.04 cm-1 (2) 

These are simply two simultaneous equations to solve. Taking 2 × (1) - (2) gives: 

 2 × [ωe – 2ωexe] - [2ωe – 6ωexe] = (2 × 2143.26 – 4260.04) cm-1

 2 ωexe = 26.48 cm-1 or ωexe = 13.24 cm-1

Substituting this value into (1) gives: 

ωe – (26.48 cm-1) = 2143.26 cm-1

 or ωe = 2169.74 cm-1